OpenStudy (anonymous):

change each fraction to a percent. 9/10, 1/16, 5/12, and 5/4

OpenStudy (anonymous):

i mean for the last one its. 4/5

OpenStudy (anonymous):

just divide them out.. and multiply by 100 again to get your percents

OpenStudy (anonymous):

9/10= 90

OpenStudy (anonymous):

1/16=6.25%

OpenStudy (anonymous):

5/12=41.67%

OpenStudy (anonymous):

4/5=80%